Ronaldo Inc. has a capital budget of $1,000,000, but it wants to maintain a target capital structure of 20% debt and 80% equity. The company forecasts this year’s net income to be $1,000,000. If the company follows a residual dividend policy, what will be its dividend payout ratio?

a.

30%

b.

20%

c.

40%

d.

50%
Answer #1

dividend payout=Net income-(Capital budget*Weight of equity)

=1,000,000-(1,000,000*80%)=$200,000

dividend payout ratio=dividend payout/net income

=200,000/1,000,000

=20%